OpenStudy (anonymous):

sin squared of x + sin x = 0 ? Find all solutions to the equation.

OpenStudy (mathmate):

\(sin^2(x)+sin(x)=0\) \(sin(x)(sin(x)+1)=0\) => sin(x)=0, sin(x)=-1 Remember "all solutions" means you have to add 2k\(\pi\) to your solution between 0 and 2\(\pi\), where k is an integer. Can you find the solutions between 0 and 2\(\pi\)?